Summary of 3mm GMVA Session October 11-16, 2007 ------------------------------------------------- This session was probably sucessfull, all stations particpated. Pico Veleta and Metsahovi are affected for a fraction of their time by bad weather. At EB and PdB the weather was very good, no major losses. In USA, VLBA_NL was partly affected by bad weather, for the other stations the conditions were moderate to good. Here some details: Pico: ------- Day 284-285 Good weather conditions.NRAO150,OJ287 and BLLAC observed normally. Day 286 Good weather. BLLAC, 0528+134, 0716+714 and OJ287 ok. Day 287 Good weather for 3C454,0528+134 and poor or unstable weather for 1803+784 Day 288 Good weather in the night for 3C84 and not very good for 3C345,3c111 and 1749+096. 24 scans lost. Day 289 Bad weather in the night 23 scans lost.16 good scans on 3c111 PdB ---- no major problems reported. Observing for a fraction of time with 5 out of 6 antennas. Onsala: ------- We did not have any major technical problems. We had a few problems with the encoder (see log for details). Howevever, as far as I know we did not loose any data. The weather has been OK (for being October in Sweden). No severe rain etc. Again, I *really* would like to have a short ftp-test before the session so I know we are OK. Metsahovi: ---------- We could run the whole schedule without any apparent problems, and both LHC+RHC seemed to stay alive in the receiver throughout the session. The weather was actually pretty good from Friday evening until Sunday afternoon; the start of c072a as well as Monday were plagued by rain. 2007.285.04:57:31.57;"Weather has deteriorated badly, raining all the time... 2007.285.13:25:03.91;"Weather has improved again, pretty clear. 2007.286.05:39:16.20;"Weather still pretty good. 2007.286.11:58:22.47;"Weather still very good, no clouds. 2007.287.08:09:12.05;"Weather is still reasonable, though quite cloudy. Rain expected later... 2007.287.12:45:28.51;"Weather: raining quite heavily already an hour, going to rain several hours at least. 2007.287.14:51:52.08;"Weather still very cloudy, though rain has ceased. 2007.288.09:54:30.09;"Weather still very cloudy, some rain occasionally and over last night, too. 2007.288.14:44:44.27;"Dull cloudy and foggy weather continues. 2007.288.17:35:27.19;"Weather cloudy, but no rain, nor predicted until end of experiment on Tue. 2007.289.04:00:15.78;"Weather somewhat cloudy. Effelsberg: ----------- Very good session, most of the time clear sky. Tsys=150-200 K. New sub-reflector and Hexapod drive required adjustment of elevation dependent focus curves and pointing model. Day 284 11.10 clear (NRAO150), no phasecal Day 285 12.10 overcast (BLLAC) some pointing problems, phasecal is now on Day 286 13.10 thin clouds, later clear (BLLAC) Day 287 14.10 clear (0716+71, 1803+78) pointing difficult due to source weakness Day 288 15.10 clear (3C84) Day 289 16.10 clear (3C120) Only few scans lost due to minor problems. VLBA: ------- VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C072A Observer: Krichbaum, T. Project type: GEO Obs filename: c072acrd.* Date/Day: 2007OCT11/284 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= 2327 Begin file, MK joins later VCLOCK check ok. NL at -0.104 UT1-UTC check ok (-0.20) GPS delta < 10us (LA: -7.7us) Operator: T. Perreault 2359-0001 *PT Formatter errors 21368 100 2 Header control module 2 error HC2 FIFO error FIFO 0 1PPS detected Cleared error condition 0000 Date is 2007OCT12/285. (First UT midnight of project) 0309-0314 %KP Formatter errors - reference pointing scan FMTR Header control module 2 error HC2 FIFO error Backplane 1PPS marker detected FIFO 3 1PPS detected FIFO 2 1PPS detected FIFO 1 1PPS detected Cleared on source change 0423 %PT Formatter errors - reference pointing scan 21368 There was also a Mark5 communication lost error for this site - the formatter errors and the communication lost error may be related in an overall communications problem. When PT changes from 7mm reference pointing scans using only BBCs 1-4 to 3mm scans using all 8 BBCs, the formatter errors go away. 0457 MK joins 0452-0459 %PT Formatter errors - reference pointing scan 21368 Cleared on source change 0529-0531 *FD autostowed. Power glitch. Repointed. COM 100 2 0547-0552 *PT autostowed. Power glitch. Repointed. COM 100 5 0554 New operator is Betty Ragan. 0624 MK high winds. 0722-0724 *LA focus rotation fault. Needed master clear. FRM 100 2 0743-0744 %PT formatter errors. They continue to happen 21368 on 7mm reference pointing scans (although not on all of them) and to clear when the source changes. The error always seems to be on HC2 only. 0808-0813 %PT formatter errors. 21368 0907-0910 *FD autostowed. Power glitch. COM 100 3 1157-1159 *KP autostowed. Power glitch. COM 100 2 1345 New Operator is Alan Kerr 1345 %PT FIFO errors continue. not logging each one. Only logging ones that occur during write scans. 1456-1508 *PT techs working on antenna, 3mm receiver cryo is 21458 100 12 cycling. 1551-2001 *PT taken out of observation. 21458 100 250 2001 %PT FIFO errors continue. 2001-1100 *PT 3mm cooling. 21458 50 449 2145 New operator: T. Perreault 2233 KP wea: clear, no precipitation 2241-2242 *PT Mark5 recording off so recorder tech can work on 21368 100 1 the formatter. 2250-2300 *PT removed from observing - formatter clock is off 21368 100 10 and incrementing 3 minutes every second. 2304-2323 *PT site tech working on formatter, no recording. 21368 100 19 2001-2323 *All data since 2001 UT is bad. Module MET-0003 21368 100 202 has 189GB of bad data. When the Mark5 comes back online, will erase this pack and start fresh. 2323 *PT formatter errors have been fixed. Pack MET-0003 has been erased, observing has resumed. 2332-1523 %PT formatter error messages continuing to appear. 21368 3 158 They are happening on both the 7mm and 3mm scans but are intermittent, no discernable pattern to it. Walter Brisken has looked into it and he believes the error messages are meaningless, not a real problem. Will count a small amount of downtime for each occurance in this one general entry instead of keeping track of each one, clearing the errors does cause a brief interruption in the data. 0000 Date is 2007OCT13/286. 0530 %OV wea: overcast, intermittent light rain 0545 New operator is Betty Ragan. 0603 *PT air handler alarm is going off, building 21464 has gotten warm. Nelson is going out to deal with it 1344 New operator is Heidi Gerhardt 1743 %FD 35+ mph wind caution 1748 %OV 35+ mph wind caution 2150 New operator: T. Perreault 0000 Date is 2007OCT14/287. 0145 New Operator: Paul Padilla 0637 %FD high winds 0946-0947 *PT autostowed, power glitch. COM 100 1 1247-1342 *NL ARACK Power supplies seem to have died, 21467 100 55 calling out Site Tech to check on it. 1342-1709 *NL 3mm running a little warm since the A rack went 21467 50 103 out. 1345 New operator is Heidi Gerhardt 1352 *NL rainging 1709-0800 *NL taken out due to 3mm warming from earlier 21467 100 891 problem with the A rack. Site tech going out to warm up 3mm in order to fix it. 0000 Date is 2007OCT15/288. 0143 New Operator: Paul Padilla 0755 %NL 3mm Cryo still warm, but is cooling enough to 21467 put back into the observation. TSys may still be off for the next hour or two. 0808-0812 *OV Formatter error on HC2 FMTR 100 4 1223 FD Site Techs report clear and dry conditions 1344 New operator is Jim Ogle. 1344-1345 *NL rotation position error. FRM 100 1 1359 LA reports clear skies. 1407-1408 *LA rotation position error. FRM 100 1 1434 *Raining at NL. 1522 KP reports clear skies. 1552-1553 *FD rotation position error. FRM 100 1 1640-1648 *NL site tech needs on antenna for a few minutes to 20467 100 8 check on the water leak problem that shorted out the A rack earlier in the observation. 1707-1709 *NL rotation position error. FRM 100 2 1707-1709 *FD rotation position error. FRM 100 2 1927 MK reports overcast skies. 2023-2025 *FD and LA rotation position errors. FRM 200 4 2101 KP reports clear skies. 2141-2142 *PT autostowed, commercial power glitch. Repointed. COM 100 1 2148 New Operator is Alan Kerr 0000 Date is 2007OCT16/289. 0055 %NL precipitation starting. 0138 NL precipitation stopped. 0220 %NL precipitation starting. 0230-0255 %OV lost communications. 0255-0855 %OV on modem. 0524 New operator: T. Perreault 1322 FD wea: clear, no precipitation 1322-1338 *NL subreflector drifting in ROTATION; ROTATION FRM 100 16 brake is not engaging allowing subreflector to drift off position at both 7mm and 3mm. 1345 New operator is Jim Ogle. 1523 End. Downtime Summary: Total downtime : 2207 min Percentage downtime of observing: 4.1% Average downtime per hour : 2.5 min Total scheduled observing time (# Antennas): 53728 min (8) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.